OpenStudy (iwanttogotostanford):

HELP PLEASE PLEASE:

myininaya (myininaya):

do you remember how to multiply conjugates for example how do you multiply (a-b)(a+b) ?

OpenStudy (iwanttogotostanford):

@myininaya no, could I have help with that?

OpenStudy (ineedhelplz):

\[√(1-\sinθ)(1+\sinθ) = √(1-\sin^2θ) = √(\cos^2θ) = \cos θ \]

OpenStudy (iwanttogotostanford):

@ineedhelplz is that the answer?